# General Compression Framework for Efficient Transformer Object Tracking

### v1

@@ Abstract
Transformer-based trackers have established a dominant role in the field of visual object tracking. While these trackers exhibit promising performance, their deployment on resource-constrained devices remains challenging due to inefficiencies. To improve the inference efficiency and reduce the computation cost, prior approaches have aimed to either design lightweight trackers or distill knowledge from larger teacher models into more compact student trackers. However, these solutions often sacrifice accuracy for speed. Thus, we propose a general model compression framework for efficient transformer object tracking, named CompressTracker, to reduce the size of a pre-trained tracking model into a lightweight tracker with minimal performance degradation. Our approach features a novel stage division strategy that segments the transformer layers of the teacher model into distinct stages, enabling the student model to emulate each corresponding teacher stage more effectively. Additionally, we also design a unique replacement training technique that involves randomly substituting specific stages in the student model with those from the teacher model, as opposed to training the student model in isolation. Replacement training enhances the student model’s ability to replicate the teacher model’s behavior. To further forcing student model to emulate teacher model, we incorporate prediction guidance and stage-wise feature mimicking to provide additional supervision during the teacher model’s compression process. Our framework CompressTracker is structurally agnostic, making it compatible with any transformer architecture. We conduct a series of experiment to verify the effectiveness and generalizability of CompressTracker. Our CompressTracker-4 with 4 transformer layers, which is compressed from OSTrack, retains about $\mathbf{96\%}$ performance on LaSOT ($\mathbf{66.1\%}$ AUC) while achieves $\mathbf{2.17\times}$ speed up.
@@ 1 Introduction
To reduce the inference cost of models, previous works attempt to design lightweight trackers or transfer the knowledge from teacher models to student trackers. Despite achieving increased speed, these existing methods still exhibit notable limitations. (1) Inferior Accuracy. Certain works propose lightweight tracking models [6, 10, 4, 22, 28] or employ neural architecture search (NAS) to search better architecture [45]. Due to the limited number of parameters, these models often suffer from underfitting and inferior performance. (2) Complex Training. Some works [15] aim to enhance the accuracy of fast trackers through transferring the knowledge from a teacher tracker to a student model. Despite the improved performance,  [15] introduces a complex multi-stage training strategy, which is time-consuming. Any suboptimal performance in these individual stages can cumulatively result in suboptimal performance in the final model. (3) Stucture Limitation. Additionally, the model reduction paradigm in [15] severely restricts the structure of student models to be consistent only with the teacher’s model.
@@
Thus, we introduce CompressTracker, a novel and general model compression framework to enhance the efficiency of transformer tracking models. The current dominant trackers are one-stream models [49, 15, 4, 10] characterized by a series of sequential transformer encoder layers, each designed to refine the temporal matching features across frames. The output of each layer is a critical temporal matching result that is refined as the layers get deeper. Given this layer-wise refinement, it becomes a natural progression to consider the model not as a single entity but as a series of interconnected stages and encourage student tracker to align teacher model at each stage. We propose the stage division strategy, which involves partitioning the teacher model, a complex pretrained transformer-based tracking model, into distinct stages that correspond to the layers of a simpler student model. This is achieved by dividing the teacher model into a number of stages equivalent to the student model’s layers. Each stage in the student model is then tasked with learning and replicating the functional behavior of its corresponding stage in the teacher model. This division is not merely a structural alteration but a strategic educational approach. By focusing each stage of the student model on mimicking a specific stage of the teacher, we enable a targeted and efficient transfer of knowledge. The student model learns not just the ’what’ of tracking—i.e., the raw matching of features—but also the ’how’—i.e., the strategies developed by the teacher model at each layer of processing.
@@
Contrary to conventional practices that isolate the training of student models, we employ a replacement training methodology that strategically intertwines the teacher and student models. The core of this methodology is the dynamic substitution of stages during training. we randomly select stages from the student model and replace them with the corresponding stages from the teacher model. By doing so, we situate the teacher model and the student model within a collaborative environment. This arrangement permits the unaltered stages of the teacher model to collaboratively inform and enhance the learning of the substituted stages in the student model rather than supervising the entire student model as a single entity. The student model is not merely learning in parallel but is directly engaging with the teacher’s learned behaviors. After training, we can just combine each stage of student model for inference. The replacement training leads to a more authentic replication of the teacher’s tracking strategies and helps to prevent the student model from overfitting to specific stages of the teacher model, promoting a more stable training.
@@
To augment the learning process, we introduce prediction guidance, which serves as a supervisory signal for the student model by leveraging the teacher model’s predictions. By using the predictions of the teacher model as a reference, the student model can converge more quickly. Furthermore, to enhance the similarity of the temporal matching features across corresponding stages, we have developed a stage-wise feature mimicking strategy. This approach systematically aligns the feature representations learned at each stage of the student model with those of the teacher model, thereby promoting a more accurate and consistent learning. In Figure 1 (a), we show the procedure and the results we are able to achieve with each step toward an efficient transformer tracker.
Figure 2: CompressTracker Framework. (a) In the training phase, we divide both the teacher model and student model into an identical number of stages. We implement a series of training strategies including replacement training, prediction guidance, and stage-wise feature mimicking, to enhance the student model’s ability to emulate the teacher model. The dotted lines represent the randomly selected paths for replacement training, with black dotted lines indicating the chosen path, while grey dotted lines denote paths not selected in a specific training iteration. (b) During inference process, we simply combine each stage of the student model for testing purposes.
@@
Compared to previous works, our CompressTracker holds many merits. (1) Enhanced Mimicking and Performance. CompressTracker enables the student model to better mimic the teacher model, resulting in better performance. As shown in Figure 1, our CompressTracker-4 achieves $\mathbf{2.17\times}$ speed up while maintaining about $\mathbf{96\%}$ accuracy. (2) Simplified Training Process. Our CompressTracker streamlines training into a single but efficient step. This simplification not only reduces the time and resources required for training but also minimizes the potential for sub-optimal performance associated with complex procedures. The training process for CompressTracker-4 requires merely 20 hours on 8 NVIDIA RTX 3090 GPUs. (3) Heterogeneous Model Compression. Our stage division strategy gives a high degree of flexibility in the design of the student model. Our framework supports any transformer architecture for student model, which is not restricted to the same structure of teacher tracker. The number of layers and their structure are not predetermined but can be tailored to fit the specific computational constraints and requirements of the deployment environment.
@@
Our contribution can be summarized as follows: (1) We introduce a novel and general model compression framework, CompressTracker, to facilitate the efficient transformer-based object tracking. (2) We propose a stage division strategy that enables a fine-grained imitation of the teacher model at the stage level, enhancing the precision and efficiency of knowledge transfer. (3) We propose the replacement training to improve the student model’s capacity to replicate the teacher model’s behavior. (4) We further incorporate the prediction guidance and feature mimicking to accelerate and refine the learning process of the student model. (5) Our CompressTracker breaks structural limitations, adapting to various transformer architectures for student model. It outperforms existing models, notably accelerating OSTrack [49] by $\mathbf{2.17\times}$ while preserving approximately $\mathbf{96\%}$ accuracy.
@@ 2 Related Work
Visual Object Tracking. Visual object tracking aims to localize the target object of each frame based on its initial appearance. Previous tracking methods [2, 30, 51, 46, 47, 16, 29, 5, 24, 12, 44] utilize a two-stream pipeline to decouple the feature extraction and relation modeling. Recently, the one-stream pipeline hold the dominant role.  [49, 14, 15, 1, 40, 25, 31, 8, 11, 20, 52] combine feature extraction and relation modeling into a unified process. These models are built upon vision transformer, which consists of a series of transformer encoder layers. Thanks to a more adequate relationship modeling between template and search frame, one-stream models achieve impressive performance. However, these models suffer from low inference efficiency, which is the main obstacle to practical deployment.
@@
Transformer Compression. Model compression aims to reduce the size and computational cost of a large model while retaining as much performance as possible, and recently many attempts have been made to speed up a large pretrained tranformer model.  [19] reduced the number of parameters through pruning technique, and [38] accomplished the quantization of BERT to 2-bits utilizing Hessian information. [37, 39, 27, 41] leverage the knowledge distillation to transfer the knowledge from teacher to student model and exploit pretrained model. Beyond language models, considerable focus has also been placed on compressing vision transformer models.  [36, 43, 9, 18, 21, 7, 48, 50] utilize multiple model compression techniques to compress vision transformer models. MixFormerV2 [15] proposed a two-stage model reduction paradigm to distill a lightweight tracker, relying on the complex multi-stage distillation training. However, our CompressTracker propose an end-to-end and efficient compression training to achieve any transformer structure compression, which speed up OSTrack $\mathbf{2.17\times}$ while maintaining about $\mathbf{96\%}$ accuracy.
@@ 3.1 Stage Division
Recently, transformer-based one-stream tracking models [8, 14, 49, 15] surpass conventional Siamese trackers [2, 13, 12], becoming the dominant manner in the field of visual object tracking. These models consist of several transformer encoder layers, each generating and progressively refining temporal matching features. Building upon this layer-wise refinement, we introduce the stage division strategy, which segments the model into a series of sequential stages. This approach encourages the student model to emulate the teacher model’s behavior at each individual stage. Specifically, we denote the pretrained tracker and the compressed model as teacher and student model, with $N_{t}$ and $N_{s}$ layers, respectively. Both teacher and student models are then divided into $N_{s}$ stages, where each stage in the student model encompasses a single layer, and each corresponding stage in the teacher model may aggregate multiple layers. For a specific stage $i$, we establish a correspondence between the stages of the teacher and student models. The objective of stage division is to enforce each stage of the student model to replicate its counterpart in the teacher model. This stage division strategy breaks the traditional approach that treats the model as an indivisible whole [6, 10, 4, 15]. Instead, it enables a fine-grained learning process where the student model transfers knowledge from the teacher in a more detailed, stage-specific manner.
@@
Unlike the reduction paradigm adopted in [15], which confines itself to pruning within identical structures, our CompressTracker framework facilitates support for arbitrary transformer structures of the student tracker, thanks to our innovative stage-wise division design. To align the size and channel dimensions of the student model’s temporal matching features with those of the teacher model, we implement input and output projection layers before and after the student layers, respectively. These projection layers serve as an adjustment mechanism to ensure compatibility between the teacher and student models and allow for a broader range of architectural possibilities for the student model. During the inference process, these input and output injection layers are omitted.
@@ 3.2 Replacement Training
During the training process, we adopt the replacement training to integrates teacher model and student models, diverging from the conventional practice of training the student model in isolation. In a specific training iteration, we implement a stochastic process to determine which stages of the student model are to be replaced by the corresponding stages of the teacher model. For the specific stage $i$, we decide whether to replace or not by random Bernoulli sampling $b_{i}$ with probability $p$, where $b_{i}\in\{0,1\}$. If $b_{i}$ equals $1$, the output from the preceding stage $i-1$ is directed to the $i$ student stage, otherwise, we channel the output into the $i$ frozen teacher stage. This replacement training creates a collaborative learning environment where the teacher model dynamically supervises the student model. The unreplaced stages of teacher provide valuable contextual supervision for a specific stage in the student model. Consequently, the student model is not operating in parallel but is actively engaged with and learning from the teacher’s established behaviors. For the optimization of student model, we only require the groundtruth box and denote the loss as $L_{track}$. Upon completion of the training process, the student model’s stages are harmoniously combined for inference. We show the pseudocode code in Appendix A.1.

### v2

@@ Abstract
Previous works have attempted to improve tracking efficiency through lightweight architecture design or knowledge distillation from teacher models to compact student trackers. However, these solutions often sacrifice accuracy for speed to a great extent, and also have the problems of complex training process and structural limitations. Thus, we propose a general model compression framework for efficient transformer object tracking, named CompressTracker, to reduce model size while preserving tracking accuracy. Our approach features a novel stage division strategy that segments the transformer layers of the teacher model into distinct stages to break the limitation of model structure. Additionally, we also design a unique replacement training technique that randomly substitutes specific stages in the student model with those from the teacher model, as opposed to training the student model in isolation. Replacement training enhances the student model’s ability to replicate the teacher model’s behavior and simplifies the training process. To further forcing student model to emulate teacher model, we incorporate prediction guidance and stage-wise feature mimicking to provide additional supervision during the teacher model’s compression process. CompressTracker is structurally agnostic, making it compatible with any transformer architecture. We conduct a series of experiment to verify the effectiveness and generalizability of our CompressTracker. Our CompressTracker-SUTrack, compressed from SUTrack, retains about $\mathbf{99\%}$ performance on LaSOT ($\mathbf{72.2\%}$ AUC) while achieves $\mathbf{2.42\times}$ speed up. Code is available at here.
@@ 1 Introduction
Previous works attempt to reduce model inference cost through lightweight tracker design or knowledge transfer from teacher to student models. However, these approaches face several limitations. (1) Inferior Accuracy. Certain works propose lightweight tracking models [6, 11, 4, 21, 31] or employ neural architecture search (NAS) [47] often suffer from underfitting and inferior performance because of the limited number of parameters. (2) Complex Training. Some works [15] transfers the knowledge from a teacher tracker to a student model, while effective, requiring complex multi-stage and time-consuming training procedures and resulting in suboptimal performance. (3) Structure Limitation. The model reduction paradigm in [15] severely restricts the structure of student to be consistent with teacher.
@@
Thus, we introduce CompressTracker, a novel and general model compression framework to enhance the efficiency of transformer tracking models, which is built upon three interconnected innovations. First, to address the structure limitation, we introduce the stage division strategy. Current dominant trackers are one-stream models [49, 15, 4, 11] with sequential transformer encoder layers that refine temporal matching features. Given this layer-wise refinement, it becomes a natural progression to consider the model not as a single entity but as a series of interconnected stages and encourage student tracker to align teacher model at each stage. Stage division partitions the teacher model into distinct stages that correspond to the layers of a simpler student model. This is achieved by dividing the teacher model into a number of stages equivalent to the student model’s layers. Each stage in the student modellearns to replicate the functional behavior of its corresponding teacher stage. This division is not merely a structural alteration but a strategic educational approach, which allows student to grasp not only what to track but also how, facilitating a more effective and fine-grained compression process.
@@
Secondly, the stage division lays the foundation for our replacement training, which tackles the complex training challenge by dynamically integrating teacher and student models during training. Replacement training strategically intertwines the teacher and student models. The core of this methodology is the dynamic substitution of stages during training. we randomly select stages from the student model and replace them with the corresponding stages from the teacher model. Replacement training permits the unaltered stages of the teacher model to collaboratively inform and enhance the learning of the substituted stages in the student model rather than supervising the entire student model as a single entity. The student model is not merely learning in parallel but is directly engaging with the teacher’s learned behaviors. After training, we can just combine each stage of student model for inference. The replacement training leads to a more authentic replication of the teacher’s tracking strategies and helps to prevent the student model from overfitting to specific stages of the teacher model.
@@
Thirdly, to further improve accuracy, we introduce prediction guidance and stage-wise feature mimicking. By using the predictions of the teacher model as a reference, the student model can converge more quickly. Feature mimicking systematically aligns the feature representations learned at each stage of the student model with those of the teacher model, thereby promoting a more accurate and consistent learning. Each of these three innovations is sequentially built upon the previous one, forming a cohesive and highly effective compression framework that maximizes efficiency while preserving performance. In Figure 1 (a), we show procedure and results we are able to achieve with each step toward an efficient transformer tracker.
@@
Compared to previous works, our CompressTracker holds many merits. (1) Enhanced Mimicking and Performance. CompressTracker enables the student model to better mimic the teacher model, resulting in better performance. As shown in Figure 1, our CompressTracker achieves an optimal balance between accuracy and inference speed. (2) Simplified Training Process. Our CompressTracker streamlines training into a single but efficient step. This simplification not only reduces the time and resources required for training but also minimizes the potential for sub-optimal performance associated with complex procedures. The training process for CompressTracker-4 requires merely 20 hours on 8 NVIDIA RTX 3090 GPUs. (3) Heterogeneous Model Compression. Our stage division strategy gives a high degree of flexibility in the design of the student model. Our framework supports any transformer architecture for student model, which is not restricted to the same structure of teacher tracker. The number of layers and their structure are not predetermined but can be tailored to fit the specific computational constraints and requirements of the deployment environment.
Figure 2: CompressTracker Framework. (a) In the training phase, we divide both the teacher model and student model into an identical number of stages. We implement a series of training strategies including replacement training, prediction guidance, and stage-wise feature mimicking, to enhance the student model’s ability to emulate the teacher model. The dotted lines represent the randomly selected paths for replacement training, with black dotted lines indicating the chosen path, while gray dotted lines denote paths not selected in a specific training iteration. (b) During inference process, we simply combine each stage of the student model for testing purposes.
@@
Our contribution can be summarized as follows: (1) We introduce a novel and general model compression framework, CompressTracker, to facilitate the efficient transformer-based object tracking. (2) We propose a stage division strategy that enables a fine-grained imitation of the teacher model at the stage level, enhancing the precision and efficiency of knowledge transfer. (3) We propose the replacement training to improve the student model’s capacity to replicate the teacher model’s behavior. (4) We further incorporate the prediction guidance and feature mimicking to accelerate and refine the learning process of the student model. (5) Our CompressTracker breaks structural limitations, adapting to various transformer architectures for student model. Our CompressTracker-SUTrack outperforms existing models, notably accelerating SUTrack [49] by $\mathbf{2.42\times}$ while preserving approximately $\mathbf{99\%}$ accuracy ($\mathbf{72.2}\%$ AUC on LaSOT).
@@ 2 Related Work
Visual Object Tracking. Visual object tracking aims to localize the target object of each frame based on its initial appearance. Previous tracking methods [2, 32, 51, 16, 33, 5, 23, 10, 46, 27, 25] utilize a two-stream pipeline to decouple the feature extraction and relation modeling. Recently, the one-stream pipeline hold the dominant role.  [49, 14, 15, 1, 42, 8, 12, 24, 19, 53, 28, 26, 54] combine feature extraction and relation modeling into a unified process. These models are built upon vision transformer, which consists of a series of transformer encoder layers. Thanks to a more adequate relationship modeling between template and search frame, one-stream models achieve impressive performance. However, these models suffer from low inference efficiency, which is the main obstacle to practical deployment.
@@
Transformer Compression. Model compression aims to reduce the size and computational cost of a large model while retaining as much performance as possible, and recently many attempts have been made to speed up a large pretrained transformer model.  [18] reduced the number of parameters through pruning technique, and [40] accomplished the quantization of BERT to 2-bits utilizing Hessian information. [39, 41, 30, 43] leverage the knowledge distillation to transfer the knowledge from teacher to student model and exploit pretrained model. Beyond language models, considerable focus has also been placed on compressing vision transformer models.  [38, 45, 9, 20, 7, 48, 50] utilize multiple model compression techniques to compress vision transformer models. MixFormerV2 [15] proposed a two-stage model reduction paradigm to distill a lightweight tracker, relying on the complex multi-stage distillation training. However, our CompressTracker propose an end-to-end and efficient compression training to achieve any transformer structure compression, which speed up SUTrack $\mathbf{2.42\times}$ while maintaining about $\mathbf{99\%}$ accuracy.
@@ 3.1 Stage Division
Recently, transformer-based one-stream tracking models [8, 14, 49, 15] have dominated visual object tracking, consisting of multiple transformer encoder layers that progressively refine temporal matching features. Building upon this layer-wise refinement, we introduce the stage division strategy, which segments the model into a series of sequential stages. This approach encourages the student model to emulate the teacher model’s behavior at each individual stage. Specifically, we denote the pretrained tracker and the compressed model as teacher and student model, with $N_{t}$ and $N$ layers, respectively. Both teacher and student models are then divided into $N$ stages, where each student stage encompasses a single layer, and each corresponding stage in teacher may aggregate multiple layers, which can be formulated as:
$teacher$ =$=\{stage_{1}^{t},stage_{2}^{t},...,stage_{N}^{t}\},$ (1)
$student$ =$=\{stage_{1}^{s},stage_{2}^{s},...,stage_{N}^{s}\},$ (2)
where $stage_{i}^{t}$ and $stage_{i}^{s}$ denote the corresponding stage $i$ in teacher and student model, respectively. For a specific stage $i$, we establish a correspondence between the stages of the teacher and student models. The objective of stage division is to enforce each stage of the student model to replicate its counterpart in the teacher model. This stage division strategy breaks the traditional approach that treats the model as an indivisible whole [6, 11, 4, 15]. Instead, it enables a fine-grained learning process where the student model transfers knowledge from the teacher in a more detailed, stage-specific manner.
@@
Unlike the reduction paradigm adopted in [15], our CompressTracker framework facilitates support for arbitrary transformer structures of the student tracker, thanks to our innovative stage-wise division design. To align the size and channel dimensions of the student model’s temporal matching features with those of the teacher model, we implement input and output projection layers before and after the student layers, respectively. These layers ensure compatibility between teacher and student models, allowing greater flexibility in the student model’s architecture. During inference, these projection layers are omitted.
Method LaSOT LaSOText TNL2K TrackingNet UAV123 FPS
AUC PNorm P AUC P AUC P AUC PNorm P AUC P GPU CPU
CompressTracker-OSTrack[49] 66.1 96% 75.2 70.6 45.7 96% 50.8 53.6 99% 52.5 82.1 99% 87.6 80.1 67.4 99% 88.0 228 2.17× 18 2.11×
CompressTracker-SeqTrack [12] 68.1 95% 77.3 73.4 47.9 96% 56.4 54.5 99% 54.2 83.1 98% 88.1 81.7 68.4 98% 90.1 78 1.95× 10 2.46×
CompressTracker-ODTrack [52] 70.5 96% 79.5 76.9 50.9 97% 60.8 58.2 99% 58.9 82.8 97% 87.9 81.4 69.2 98% 90.8 87 2.71× 13 3.18×
CompressTracker-SUTrack [13] 72.2 99% 80.7 78.2 52.1 98% 63.2 64.4 99% 69.9 84.3 99% 88.2 83.3 71.0 99% 91.4 134 2.42× 24 2.34×
SUTrack-T [13] 69.6 79.3 75.4 50.2 57.0 - - 82.7 87.2 80.8 69.4 - 100 23
HiT-Base [31] 64.6 73.3 68.1 44.1 - - - 80.0 84.4 77.3 65.6 - 175 33
HiT-Samll [31] 60.5 68.3 61.5 40.4 - - - 77.7 81.9 73.1 63.3 - 192 72
HiT-Tiny [31] 54.8 60.5 52.9 35.8 - - - 74.6 78.1 68.8 53.2 - 204 76
SMAT [21] 61.7 71.1 64.6 - - - - 78.6 84.2 75.6 64.3 83.9 158 33
MixFormerV2-S [15] 60.6 69.9 60.4 43.6 46.2 48.3 43.0 75.8 81.1 70.4 65.8 86.8 325 30
FEAR-L [6] 57.9 68.6 60.9 - - - - - - - - - - -
FEAR-XS [6] 53.5 64.1 54.5 - - - - - - - - - 80 60
HCAT [11] 59.0 68.3 60.5 - - - - 76.6 82.6 72.9 63.6 - 195 45
E.T.Track [4] 59.1 - - - - - - 74.5 80.3 70.6 62.3 - 150 47
LightTrack-LargeA [47] 55.5 - 56.1 - - - - 73.6 78.8 70.0 - - - -
LightTrack-Mobile [47] 53.8 - 53.7 - - - - 72.5 77.9 69.5 - - 120 41
STARK-Lightning [46] 58.6 69.0 57.9 - - - - - - - - - 200 -
DiMP [3] 56.9 65.0 56.7 - - - - 74.0 80.1 68.7 65.4 - 77 -
SiamFC++ [44] 54.4 62.3 54.7 - - - - 75.4 80.0 70.5 - - 90 -
Table 1: Compress Object Tracker. We compress several trackers such as SUTrack [13], ODTrack [52], and so on, into students with an optimal trade-off between accuracy and efficiency. We report the performance on 5 benchmarks and calculate the performance gap in comparison to the origin teacher model. CompressTracker outperforms previous models in both speed and accuracy.
@@ 3.2 Replacement Training
During training, we adopt the replacement training to integrates teacher model and student models, diverging from the conventional practice of training the student model in isolation. In each iteration, a stochastic process determines which student model stages are replaced by corresponding teacher stages. For the specific stage $i$, the forward process in conventional isolated propagation can be described as:
$h_{i}=stage_{i}^{s}(h_{i-1}),$ (3)
where $h_{i-1}$ is the input of the $i$ student stage. However, in our replacement training, we decide whether to replace or not by random Bernoulli sampling $b_{i}$ with probability $p$, where $b_{i}\in\{0,1\}$. If $b_{i}$ equals $1$, the output from the preceding $\mathit{i-1}$-th stage is directed to the $i$ student stage, otherwise, we channel the output into the $i$ frozen teacher stage, which can be formulated as:
$h_{i}=\left\{\begin{array}[]{ll}stage_{i}^{t}(h_{i-1}),&r_{i}=0,\\ stage_{i}^{s}(h_{i-1}),&r_{i}=1,\end{array}\right.r_{i}\sim\mathrm{Bernoulli}(p).$ (4)
This replacement training creates a collaborative learning environment where the teacher model dynamically supervises the student model. The unreplaced stages of teacher provide valuable contextual supervision for a specific stage in the student model. Consequently, the student model is not operating in parallel but is actively engaged with and learning from the teacher’s established behaviors. For the optimization of student model, we only require the groundtruth box and denote the loss as $L_{track}$. After training, the student model’s stages are harmoniously combined for inference.